# Inferring fine-grained migration patterns across the United States

### v2

@@ Abstract
Fine-grained migration data illuminate important demographic, environmental, and health phenomena. However, migration datasets within the United States remain lacking: publicly available Census data are neither spatially nor temporally granular, and proprietary data have higher resolution but demographic and other biases. To address these limitations, we develop a scalable iterative-proportional-fitting based method that reconciles high-resolution but biased proprietary data with low-resolution but more reliable Census data. We apply this method to produce MIGRATE, a dataset of annual migration matrices from 2010 - 2019 that captures flows between 47.4 billion pairs of Census Block Groups — about four thousand times more granular than publicly available data. These estimates are highly correlated with external ground-truth datasets, and improve accuracy and reduce bias relative to raw proprietary data. We use MIGRATE to analyze both national and local migration patterns. Nationally, we document temporal and demographic variation in homophily, upward mobility, and moving distance: for example, we find that people are increasingly likely to move to top-income-quartile CBGs and identify racial disparities in upward mobility. We also show that MIGRATE can illuminate important local migration patterns, including out-migration in response to California wildfires, that are invisible in coarser previous datasets. We publicly release MIGRATE to provide a resource for migration research in the social, environmental, and health sciences.
@@ 1 Introduction
Fine-grained migration data, which record the number of people relocating from one geographic area to another, are essential for understanding a range of social, environmental, and health phenomena. Migration data illuminate responses to environmental disasters and climate change [1, 2, 3]; responses to economic stresses and opportunities [4, 5, 6]; patterns of social change [7]; consequences of conflicts [8]; effects of the COVID-19 pandemic [9]; and political polarization [10].
@@
But migration datasets within the United States have serious limitations. The most fine-grained freely available datasets track migration at the county level [11]. While widely used, these datasets lack sufficient spatial resolution to study a number of phenomena where previous research has revealed important variation at the sub-county level. For example, research on flood-risk-induced migration uses models at the much more granular Census block level [12], arguing that this spatial granularity is necessary due to the highly localized nature of flood risk [13]. Research on neighborhoods that are vulnerable to climate-induced eviction or displacement similarly models risk at a sub-county (Census tract) level [14]. Movers within the same county have accounted for more than half of migratory flows in the United States in all years from 2006 to 2019 [15]. Publicly available, county-level migration datasets render these patterns entirely invisible.
@@
Proprietary migration datasets offer greater granularity. Data aggregators like Infutor [16] combine many data sources — including voter files, property deeds, credit files, and phone books [17] — to attempt to infer address histories for individual-level movers. Such datasets have been widely used because they are extremely temporally and spatially fine-grained [18, 4, 19, 20, 21]. Previous work also suggests these data do contain valuable signal which correlates with external datasets, as validated by cross-referencing the data with hurricanes or public housing foreclosures [18], and with marginal population counts in the region of interest [17]. However, these proprietary datasets have three major disadvantages. First, they are not publicly available, limiting their utility to researchers. Second, they require extensive computational pipelines, and substantial computational resources, to clean and map to standardized geographical areas to facilitate subsequent analysis. Third, and most fundamentally, they combine multiple imperfect data sources using proprietary algorithms, and thus contain noise and biases. For example, Infutor and other consumer record datasets have been shown to overrepresent higher-income and majority-group populations in some settings [22].
To address the limitations of existing migration data, we create and release MIGRATE (Migration Inference for GRAnular Trend Estimation): fine-grained migration estimates that combine the strengths of both aforementioned types of data by harmonizing biased but fine-grained proprietary data with reliable but coarse Census data. To produce our estimates, we develop a method based on iterative proportional fitting (IPF), an algorithm that repeatedly scales a noisy matrix to reconcile it with constraints [23, 24, 25], to reconcile the raw Infutor migration matrix with the more reliable Census constraints. We accommodate uncertainty in address mapping and incorporate data on births, deaths, and immigration. The output of our method is a set of yearly inferred migration matrices at the Census block group (CBG) level from 2010-2019 covering the 50 states and the District of Columbia. Our matrices capture migration flows between 47.4 billion pairs of CBGs, making MIGRATE approximately 4,600 times more granular than the county-county flows publicly available on the 5-year level, and 18 million times more granular than the state-state flows publicly available on the 1-year level.
@@
We comprehensively validate MIGRATE by comparing it to external data sources. We show that our estimates correlate well with fine-grained Census and ACS data on population counts and migration flows. We show that MIGRATE improves accuracy relative to raw Infutor data: in particular, it reduces the deviation from ground truth ACS county-to-county migration estimates by 90%. We document systematic demographic biases in the Infutor data, which overcounts rural, older, white, and home-owning populations, and show that MIGRATE almost entirely eliminates these biases.
We then use MIGRATE to analyze both national and local migration. Nationally, we reveal both temporal and demographic variation in migration homophily, upward mobility, and moving distance. We find, for example, that people are increasingly likely to move to top-income-quartile CBGs, but also provide evidence of racial disparities: movers from plurality Black CBGs are less likely, and movers from plurality Asian CBGs more likely, to move to top-income-quartile CBGs, and these disparities persist even when controlling for income of origin CBG. Locally, we show that MIGRATE can illuminate important migration patterns, including dramatic increases in out-migration in response to California wildfires, that are invisible in county-level data.
Collectively, these analyses illustrate how MIGRATE reveals important migration patterns at both the national and local levels. To provide a foundation for more precise migration research in the social, environmental, and health sciences, we release MIGRATE for other researchers at migrate.tech.cornell.edu.
2 Results
2.1 Creating MIGRATE
@@
Preprocessing the raw Infutor dataset
We first preprocess the raw Infutor data into yearly migration matrices. The raw Infutor data consists of sequences of addresses for individual people. We use this information to compute the number of people moving between each pair of geographic areas. In particular, we estimate the matrices $E^{(t)}$, where entries $E^{(t)}_{ij}$ represent the number of people who reside in CBG $j$ some time during year $t$ and resided in CBG $i$ a year prior [^1] We correct populations to account for birth, deaths, and international migration. Full details in the Methods Section (§M.2).. We provide descriptive statistics for the raw Infutor data and the processed migration matrices for the 2010-2019 time period in Table 1 and additional descriptive statistics in Table M1.
@@
Table 1: Summary statistics for raw Infutor data and Census data during the period of interest (2010-2019). We define “individuals with active records” as those whose [earliest Infutor date observed, latest Infutor date observed] interval intersects a given time period. Comparing the average yearly number of active individuals in the Infutor data to the Census population (second row) shows that Infutor under-counts the population. (Table M1 provides the number of active individuals in each year from 2010-2019). Active individuals have on average between two and three addresses, translating to one or two moves during the decade.
Preprocessing Infutor requires mapping addresses to geographic areas. When the raw Infutor data contains a precise street address (90.40% of all addresses in the dataset), we map each address to a unique 2010 CBG with a state-of-the-art geocoder [26, 27]. We obtain matches for 92.18% of these addresses and confirm by hand inspection of 200 addresses that this yields highly accurate results. The remaining addresses are incomplete or imprecise (e.g., contain only ZIP codes); we probabilistically map these addresses to multiple CBGs, with the weight of each CBG proportional to the population in the CBG intersecting the ZIP code. We are able to map 99.21% of all addresses in Infutor to CBGs. We fully describe the extensive additional preprocessing steps for the raw Infutor data — including accounting for uncertainty in move date, choosing between geocoders, and prioritizing addresses for individuals with conflicting histories — in the Methods Section (§M.1).
Harmonizing the Infutor and Census datasets
@@
Having produced the raw Infutor yearly migration matrices $E^{(t)}$, we next reconcile them with more reliable but less granular Census data by rescaling selected entries of $E^{(t)}$. For example, to reconcile $E^{(t)}$ with Census county populations at time $t$, we multiply the rows corresponding to each county by a scaling factor to match the Census population in that county. We rescale to different Census datasets in sequence, rather than simultaneously, because the datasets are not totally consistent with each other.
@@
Specifically, we first rescale the entries of $E^{(t)}$ to match CBG populations; then the 1-year counts of movers and non-movers by state; then the 1-year state-to-state flows; and finally the 1-year county populations. The motivation for this ordering is that the 1-year county populations are more precisely estimated than the other datasets, and we thus prioritize matching to them: while CBG populations and state flows are estimated by the American Community Survey (ACS) and often contain large sampling errors, county populations are estimated by the Population Estimates Program (PEP) using more precise large-scale administrative datasets. We verify that incorporating each of these datasets improves our performance on the validations discussed below, and that the performance is not overly sensitive to the order in which we match to the datasets (Appendix A.2). To match each matrix $E^{(t)}$ to annual county populations, we apply an IPF-based algorithm [23, 24]: we scale blocks of rows and columns to agree with the annual county populations in years $t-1$ and $t$, respectively, alternating between scaling rows and scaling columns until the procedure converges. The resulting migration matrices constitute our MIGRATE estimates. The Methods Section (§M.3) provides additional methodological details on these scalings.
@@ 2.2 Validating MIGRATE
We conduct extensive validations of the MIGRATE estimates. We first show that MIGRATE estimates are highly concordant with Census datasets and improve agreement relative to raw Infutor data (Figure 2). We then investigate demographic biases in the raw Infutor estimates and show that MIGRATE reduces them (Figure 3).
@@
Figure 2: Validating the MIGRATE estimates. (a - c): MIGRATE estimates (y-axis) are highly correlated with Census data (x-axis), including (a) Census populations at the state, county, tract, and block group level, (b) Census flows between each pair of states and each pair of counties, and (c) state and county in-migration rates (i.e., the number of people moving into an area as a fraction of the area’s population). For in-migration rates, correlations are weighted by state or county population, and points are sized by population. (d - f) MIGRATE estimates increase agreement with Census datasets relative to raw Infutor data. We compute RMSE between (1) MIGRATE estimates and Census data and (2) Infutor data and Census data, and report the reduction in RMSE from using MIGRATE estimates. Bars show the mean reduction in RMSE across all years; error bars plot standard deviation across years. For the in-migration rates, RMSE is weighted by area population.
@@
MIGRATE estimates are highly correlated with ground-truth Census data
@@
MIGRATE estimates are well-correlated with all available Census measures (Figure 2a-b). We first assess correlations with Census datasets incorporated in our harmonization procedure, to confirm it indeed helps reconcile MIGRATE with Census data. We would expect these correlations to be high but not uniformly perfect, because the Census datasets are not totally consistent with each other. By design, MIGRATE estimates perfectly match state and county populations (because the last step in our procedure is to match to county populations). MIGRATE estimates also achieve Pearson correlation $\rho=0.997$ with 5-year Census tract populations, and $\rho=0.996$ with 5-year CBG populations [^2] To compare our one-year MIGRATE estimates to five-year ACS estimates, we average our estimates across the same periods for consistency with the ACS data. We only use ACS data products whose period completely overlaps with the 2010-2019 MIGRATE range.. MIGRATE estimates are also highly correlated with state-to-state flows ($\rho=1$); this high correlation is driven in part by variation in state populations, as well as by the fact that most people do not move. Hence, we also compute correlation with state in-migration rates ($\rho=0.987$), which is not subject to either of these effects; in Appendix A.1, we report additional validations using diagonal and off-diagonal flow entries, as well as raw in-migration, which are also highly correlated with our estimates.
Importantly, we further verify that our estimation procedure yields highly correlated estimates even with datasets which are not used in our estimation procedure. This constitutes a validation on held-out data, a standard check for statistical estimation methods [28]. Specifically, we compare our MIGRATE estimates to ACS 5-year county-to-county flows, which are not incorporated in our estimation, and find they are highly correlated with our estimates ($\rho=1$). This high correlation is driven in part by the fact that most people do not move; hence, in Appendix A.1, we compute correlations separately for county movers and county non-movers (all above $\rho=0.95$), as well as raw in-migration ($\rho=0.992$). Another contributor to the high correlation in county-county flows is variation in county size, so we separately compute correlations with county in-migration rates (i.e., the number of people moving into an area as a fraction of the area’s population) which normalize for county size. Correlations remain high (population-weighted $\rho=0.715$).
As a further held-out validation, we verify that our estimation procedure yields highly correlated estimates with each Census data source even when we remove it from the datasets used for estimation: for example, we remove CBG-level Census populations from our estimation datasets, and verify that our resulting estimates remain highly correlated with CBG-level data ($\rho=0.888$ with Census Tract populations and $\rho=0.856$ with Census block group populations). See Appendix A.2 for full results.
@@
Overall, these validations demonstrate that MIGRATE estimates are highly correlated with external ground-truth Census data. This is not merely because of variation in state or county size (since it remains true when examining in-migration rates, which normalize for state or county size, as well as when looking at the tract and CBG level). Nor is it due merely to the fact that most people do not move (since it remains true when examining correlations for movers specifically). Finally, it is not merely a consequence of overfitting to Census data, because it remains true on held-out datasets that are not used in our estimation procedure.
MIGRATE estimates reduce error relative to raw Infutor data

### v3

@@ Abstract
Fine-grained migration data illuminate demographic, environmental, and health phenomena. However, United States migration data have serious drawbacks: public data lack spatiotemporal granularity, and higher-resolution proprietary data suffer from multiple biases. To address this, we develop a method that fuses high-resolution proprietary data with coarse Census data to create MIGRATE: annual migration matrices capturing flows between 47.4 billion US Census Block Group pairs—approximately four thousand times the spatial resolution of current public data. Our estimates are highly correlated with external ground-truth datasets and improve accuracy relative to raw proprietary data. We use MIGRATE to analyze national and local migration patterns. Nationally, we document demographic and temporal variation in homophily, upward mobility, and moving distance—for example, rising moves into top-income-quartile block groups and racial disparities in upward mobility. Locally, MIGRATE reveals patterns such as wildfire-driven out-migration that are invisible in coarser previous data. We release MIGRATE as a resource for migration researchers.
@@ 1 Introduction
Fine-grained migration data, which record the number of people relocating from one geographic area to another, are essential for understanding a range of social, environmental, and health phenomena. Migration data illuminate responses to environmental disasters and climate change hoffmann_drought_2024, kaysen_open_2025, mcconnell_rare_2024, alexander2019impact; responses to economic stresses and opportunities diamond_effect_2020, wilkerson_warmth_2010, niva_worlds_2023, boucherie_decoupling_2025; patterns of social change depillis_for_2024; consequences of conflicts meta-migration, leasure2023nowcasting; effects of the COVID-19 pandemic ramani_how_2024; housing instability phillips_measuring_2020, gershenson2024fracking; urban-suburban migration patterns golding2020tracking, reia2022spatial; and political polarization kaysen_millions_2024.
@@
But migration datasets within the United States have serious limitations. The most fine-grained publicly available national datasets track migration at the county level county_to_county_data. While widely used, these datasets lack sufficient spatial resolution to study a number of phenomena where previous research has revealed important variation at the sub-county level. For example, research on flood-risk-induced migration uses models at the much more granular Census block level shu2023integrating, arguing that this spatial granularity is necessary due to the highly localized nature of flood risk columbia2024climatemigration. Other research on climate-induced vulnerability, eviction or displacement similarly models risk at a sub-county (Census tract) level Tedesco2021, gershenson2024fracking or even at the household level rising2022missing. Research on housing instability often models Census Block Groups freedman_living_2022 or even specific building complexes rutan_concentrated_2021. All of these applications testify to the need for highly granular data to facilitate accurate study of many important migration-related phenomena. Additionally, movers within the same county have accounted for more than half of migratory flows in the United States in all years from 2006 to 2019 kerns-damore_migration_2023. Publicly available, county-level migration datasets are too coarse to study these important migration patterns.
@@
Proprietary migration datasets offer greater granularity. Data aggregators like Infutor infutor combine many data sources — including voter files, property deeds, credit files, and phone books diamond_effects_2019, phillips2024homelessness, boar2023consumption — to attempt to infer address histories for individual-level movers. Such datasets have been widely used because they are extremely temporally and spatially fine-grained phillips_measuring_2020, diamond_effect_2020, bernstein2022contribution, asquith2019supply, qian2021effects, phillips2024homelessness. A long literature which provides more recent or granular migration estimates using non-traditional data sources, including social media and digital advertisement data, testifies to the power of new data sources meta-migration, alexander2019impact, alexander2022combining, leasure2023nowcasting, rampazzo2025assessing. Previous work also suggests that address history data do contain valuable signal which correlates with external datasets, as validated by cross-referencing the data with hurricanes or public housing foreclosures phillips_measuring_2020, and with marginal population counts in the region of interest diamond_effects_2019. However, proprietary address history datasets have three major disadvantages. First, they are not publicly available, limiting their utility to researchers. Second, they require extensive computational pipelines, and substantial computational resources, to clean and map to standardized geographical areas to facilitate subsequent analysis. Third, and most fundamentally, they combine multiple imperfect data sources using proprietary algorithms, and thus contain noise and biases. For example, Infutor and other consumer record datasets have been shown to overrepresent higher-income and majority-group populations in some settings ramiller_residential_2024.
@@
To address the limitations of existing migration data, we create and release MIGRATE (Migration Inference for GRAnular Trend Estimation): fine-grained migration estimates that combine the strengths of both aforementioned types of data by harmonizing biased but fine-grained proprietary data with reliable but coarse Census data. To produce our estimates, we develop a data fusion methodimbens_combining_1994, guan2025datafusionhighresolutionestimation based on iterative proportional fitting (IPF)IPF, chang2024inferring, chang2021mobility to reconcile the raw migration data from data aggregator Infutor with the more reliable Census constraints. The output of our method is a set of yearly inferred United States migration matrices at the Census Block Group (CBG) level from 2010-2019. Our matrices capture migration flows between 47.4 billion pairs of CBGs, making MIGRATE approximately 4,600 times more granular than the county-county flows publicly available on the 5-year level, and 18 million times more granular than the state-state flows publicly available on the 1-year level. We comprehensively validate MIGRATE by comparing to external data sources, showing that it correlates well with ground-truth population counts and migration flows, and improves accuracy and reduces demographic bias compared to raw Infutor data, which overcounts rural, older, white, and home-owning populations. We then use MIGRATE to analyze both national and local migration. Nationally, we reveal both temporal and demographic variation in migration homophily, upward mobility, and moving distance. We find, for example, that people are increasingly likely to move to top-income-quartile CBGs, but also provide evidence of racial disparities: movers from plurality Black CBGs are less likely, and movers from plurality Asian CBGs more likely, to move to top-income-quartile CBGs, and these disparities persist even when controlling for income of origin CBG. Locally, we show that MIGRATE can illuminate important migration patterns, including dramatic increases in out-migration in response to California wildfires, that are invisible in county-level data. To provide a foundation for more precise migration research in the social, environmental, urban, and health sciences, we release MIGRATE for non-profit research use at migrate.tech.cornell.edu.
2 Creating MIGRATE
@@
2.1 Preprocessing the raw Infutor dataset
We first preprocess the raw Infutor data into yearly migration matrices. The raw Infutor data consists of sequences of addresses for individual people [^1] For a subset of individuals, Infutor also provides estimated demographic information (e.g., age, gender, and socioeconomic information). We do not incorporate this information in our analysis for two reasons: first, complete demographic data is only available for about half of all individuals; second, the reliability of estimated demographic information is concerning—as the data lacks documentation, we should not discard the possibility that demographic information has been imputed based on addresses—and contains known and potential biases ramiller_residential_2024, 10.1093/pnasnexus/pgaf027. Analogously, we do not use Census information on demographic subgroups when harmonizing with Census data; we only make use of population counts and flows.. We use these address sequences to compute the number of people moving between each pair of geographic areas. In particular, we estimate the matrices $E^{(t)}$, where entries $E^{(t)}_{ij}$ represent the number of people who reside in CBG $j$ some time during year $t$ and resided in CBG $i$ a year prior.
We provide a conceptual overview of the preprocessing pipeline here and full details in the Methods Section M.1. We first construct cleaned monthly address histories for each individual in the Infutor dataset by reconciling inconsistencies in address start/end dates, discarding unreliable records (e.g., postal boxes), and modeling uncertainty when multiple addresses are active. These monthly address distributions are then aggregated in a way that simulates ACS survey responses about residence one year prior, producing annual migration flow estimates between pairs of addresses. Finally, to convert the address-address migration matrices to CBG-CBG migration matrices, each address is mapped to one or more CBGs. When the raw Infutor data contains a precise street address (90.40% of all addresses in the dataset), we map each address to a unique 2010 CBG with a state-of-the-art geocoder R-arcgeocoder, us_census_bureau_census_2024. We obtain matches for 92.18% of these addresses and confirm by hand inspection of 200 addresses that this yields highly accurate results. The remaining addresses are incomplete or imprecise (e.g., contain only ZIP codes); we probabilistically map these addresses to multiple CBGs, with the weight of each CBG proportional to the population in the CBG intersecting the ZIP code. We are able to map 99.21% of all addresses in Infutor to CBGs. The final output of this preprocessing procedure is the Infutor yearly migration matrices $E^{(t)}$, which capture migration counts between pairs of CBGs as estimated from Infutor data alone. We provide summary statistics for the raw Infutor data and the processed migration matrices for the 2010-2019 time period in Table 1 and additional descriptive statistics in Table 2.
@@ 2.1 Preprocessing the raw Infutor dataset
Table 1: Summary statistics for raw Infutor data and Census data (2010-2019). We define “individuals with active records” as those whose [earliest Infutor date observed, latest Infutor date observed] interval intersects a given time period. Comparing the average yearly number of active individuals in the Infutor data to the Census population (second row) shows that Infutor under-counts the population. (Table 2 provides the number of active individuals in each year from 2010-2019). Active individuals have on average between two and three addresses, translating to one or two moves during the decade.
2.2 Harmonizing the Infutor and Census datasets
@@
Having produced the raw Infutor yearly migration matrices $E^{(t)}$, we next reconcile them with more reliable but less granular Census data by rescaling selected entries of $E^{(t)}$. We rescale to different Census datasets in sequence, rather than simultaneously, because the datasets are not totally consistent with each other. Specifically, we first rescale the entries of $E^{(t)}$ to match CBG populations from the Census and 5-year American Community Survey (ACS); then the 1-year ACS counts of movers and non-movers by state; then the 1-year ACS state-to-state flows; and finally the 1-year county populations from the Population Estimates Program (PEP).
@@
The motivation for this ordering is that the 1-year county populations are more precisely estimated than the other datasets, and we thus prioritize matching to them. While CBG populations and state flows are estimated by the ACS and often contain large sampling errors (see our discussion in Methods Section M.2), county populations are estimated by PEP using more precise large-scale administrative datasets. We verify that incorporating each of these datasets improves our performance on the validations discussed below, and that the performance is not overly sensitive to the order in which we match to the datasets (see Supplementary Table 1).
To perform the final rescaling (matching 1-year county populations) we apply an iterative procedure based on the classical IPF algorithm IPF, chang2024inferring: specifically, we scale blocks of rows and columns to agree with the annual county populations in years t1t-1 and $t$, respectively, alternating between scaling rows and scaling columns until the procedure converges. Only the final rescaling is performed using full IPF; the other rescalings, to CBG populations, 1-year counts of movers and non-movers, and 1-year state-to-state flows, are performed only once at the beginning of the process, as opposed to iteratively, to avoid overfitting to these more noisily estimated datasets. The resulting migration matrices constitute our MIGRATE estimates. In Methods Section M.3, we provide additional details on these scalings.
@@ 3 Validating MIGRATE
We conduct extensive validations of the MIGRATE estimates. We first show that MIGRATE estimates are highly concordant with Census datasets and improve agreement relative to raw Infutor data (Figure 2). We then investigate demographic biases in the raw Infutor estimates and show that MIGRATE reduces them (Figure 3). In Appendix A.2 we further verify, with experiments on semi-synthetic data, that our method recovers ground-truth flows under certain conditions of bias and noise.
@@
Figure 2: Validating the MIGRATE estimates. (a - c): MIGRATE estimates (y-axis) are highly correlated with Census data (x-axis), including (a) Census populations at the tract, and block group (CBG) level, (b) movers between each pair of states and each pair of counties (excluding people who remain within the same state or county), and (c) state and county in-migration rates (i.e., the number of people moving into an area as a fraction of the area’s population). For in-migration rates, correlations are weighted by state or county population, and points are sized by population. (d - f) MIGRATE estimates increase agreement with Census datasets relative to raw Infutor data for population counts, movers between states and counties, and in-migration rate, respectively. We compute root mean squared error (RMSE) between (1) MIGRATE estimates and Census data and (2) Infutor data and Census data, and report the reduction in RMSE from using MIGRATE estimates. Bars show the mean reduction in RMSE across all data release years ($n=5$ for population and county-level migration datasets, $n=9$ for state-level migration datasets); error bars plot standard deviation across years. For the in-migration rates, RMSE is weighted by area population.
@@
3.1 MIGRATE estimates are highly correlated with ground-truth Census data
@@
MIGRATE estimates are well-correlated with all available Census measures of population counts and flows (Figure 2a-c). By design, MIGRATE estimates perfectly match state and county populations (because the last step in our harmonization procedure is to match to county populations). MIGRATE estimates also achieve Pearson correlation $\rho=0.997$ with 5-year Census tract populations, and $\rho=0.996$ with 5-year CBG populations (Figure 2a) [^2] To compare our one-year MIGRATE estimates to five-year ACS estimates, we average our estimates across the same periods for consistency with the ACS data, using only ACS data products whose time period completely overlaps with the 2010-2019 MIGRATE range.. MIGRATE estimates of the number of movers between each pair of states and each pair of counties are also highly correlated with ACS estimates ($\rho=0.998$ and $0.957$ for states and counties respectively; Figure 2b) [^3] We exclude people who remain within the same state or county from this calculation, because most people do not move, which artificially inflates the correlation. Supplementary Table 3(b) reports the correlation without this exclusion, which is nearly perfect.. Finally, MIGRATE estimates of state and county in-migration rates (i.e., the number of people moving into an area as a fraction of the area’s population) are well-correlated with ACS estimates (Figure 2c; $\rho=0.987$ and 0.715 for state and county in-migration rates, respectively, weighting by area population). Supplementary Table 1(a) reports correlations with additional Census quantities (e.g., in-migration counts as opposed to rates) which remain high.
@@
Overall, these validations demonstrate that MIGRATE estimates are highly correlated with ground-truth Census data. This is not merely because of variation in state or county populations, since it remains true when examining in-migration rates, which normalize for population, as well as when looking at tracts and CBGs, which do not vary as significantly in population. Nor is it due merely to the fact that most people do not move, since we examine correlations for movers specifically, as well as correlations with in-migration rates. Finally, it is not merely a consequence of overfitting to Census data, because it remains true on held-out datasets that are not used in our harmonization procedure, a standard check for statistical estimation methods hastie_elements_2009444Even for some Census datasets that are used in our harmonization procedure, we would expect correlations with MIGRATE to be high but not perfect, because the Census datasets are not totally consistent with each other.. Specifically, MIGRATE estimates remain highly correlated with county-county mover counts and county in-migration rates, which are not used in our estimation procedure. As a further held-out validation, we verify that our estimation procedure yields highly correlated estimates with each Census data source even when we remove it from the datasets used for estimation. For example, we remove CBG-level Census populations from our estimation datasets, and verify that our resulting estimates remain highly correlated with data below county level ($\rho=0.888$ with Census Tract populations and $\rho=0.856$ with Census Block Group populations). See Supplementary Table 1(a) for full results.
3.2 MIGRATE estimates reduce error relative to raw Infutor data